# Data-integrated neural networks for solving partial differential equations

### v3

@@ Abstract.
In this work, we propose data-integrated neural networks (DataInNet) for solving partial differential equations (PDEs), offering a novel approach to leveraging data (e.g., source terms, initial conditions, and boundary conditions). The core of this work lies in the integration of data into a unified network framework. DataInNet comprises two subnetworks: a data integration neural network responsible for accommodating and fusing various types of data, and a fully connected neural network dedicated to learning the residual physical information not captured by the physics integration neural network. This network architecture inherently excludes function classes that violate known physical constraints, thereby substantially narrowing the solution search space. Numerical experiments demonstrate that the proposed DataInNet delivers superior performance on challenging problems, such as the Helmholtz equation (relative $L^{2}$ error: O($10^{-6}$)) and PDEs with high frequency solutions (relative $L^{2}$ error: O($10^{-5}$)).
@@ 1. Introduction
In scientific computing, deep neural networks (DNNs) have been widely adopted for solving various partial differential equations (PDEs) [1, 2, 3, 4, 5, 6, 7]. While notable progress has been made in PDE solving using DNNs [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], several critical challenges remain unaddressed, such as spectral bias (i.e., the frequency principle) [19]. This bias is characterized by the tendency of DNNs to prioritize learning low frequency components over high frequency ones when fitting functions or solving PDEs. The existence of spectral bias has been experimentally validated in numerous studies [20, 21, 22], and Tancik et al. leveraged the neural tangent kernel [23] to provide corresponding theoretical justification. Inspired by insights into the neural tangent kernel, physics-informed neural networks incorporate multi-scale random Fourier feature mapping to enhance robustness and accuracy [24, 25]. Furthermore, PINNs leverage the derivatives of the solution to define the loss function-a design that may facilitate the convergence of high-frequency components during training. Examples presented by Lu et al. [21] confirm that differentiation promotes the convergence of high frequency components.
@@ 2.2. Data-integrated neural networks
DataInNet comprises a data integration neural network and an auxiliary neural network. The data integration neural network is specifically designed to accept and integrate various types of data, including source terms, initial conditions, and boundary conditions. The auxiliary neural network is a fully connected network tasked with learning the residual physical information not captured by the physics integration neural network. The forward propagation rule for data integration neural networks with $N^{*}$ modules are as follows:
@@ 4. Summary and discussion
Numerical experiments demonstrate the effectiveness of DataInNet. These experiments encompass the high-wavenumber Helmholtz equation, the Poisson equation defined on an L-shaped domain, and PDEs with high-frequency solutions. Future research will focus on devising more efficient data integration methods and assessing the applicability of this approach in more challenging scenarios.

### v4

@@ Abstract.
In this work, we propose data-integrated neural networks (DataInNet) for solving partial differential equations (PDEs), offering a novel approach to leveraging data (e.g., source terms, initial conditions, and boundary conditions). The core of this work lies in the integration of data into a unified network framework. DataInNet comprises two subnetworks: a data integration neural network responsible for accommodating and fusing various types of data, and a fully connected neural network dedicated to learning the residual physical information not captured by the data integration neural network. This network architecture inherently excludes function classes that violate known physical constraints, thereby substantially narrowing the solution search space. Numerical experiments demonstrate that the proposed DataInNet delivers superior performance on challenging problems, such as the Helmholtz equation (relative $L^{2}$ error: O($10^{-6}$)) and PDEs with high frequency solutions (relative $L^{2}$ error: O($10^{-5}$)).
@@ 1. Introduction
In scientific computing, deep neural networks (DNNs) have been widely adopted for solving various partial differential equations (PDEs) [1, 2, 3, 4, 5, 6, 7]. While notable progress has been made in PDE solving using DNNs [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], several critical challenges remain unaddressed, such as spectral bias (i.e., the frequency principle) [19]. This bias is characterized by the tendency of DNNs to prioritize learning low frequency components over high frequency ones when fitting functions or solving PDEs. The existence of spectral bias has been experimentally validated in numerous studies [20, 21, 22], and Tancik et al. leveraged the neural tangent kernel [23] to provide corresponding theoretical justification. Inspired by insights into the neural tangent kernel, physics-informed neural networks (PINNs) incorporate multi-scale random Fourier feature mapping to enhance robustness and accuracy [24, 25]. Furthermore, PINNs leverage the derivatives of the solution to define the loss function-a design that may facilitate the convergence of high frequency components during training. Examples presented by Lu et al. [21] confirm that differentiation promotes the convergence of high frequency components.
@@ 2.2. Data-integrated neural networks
DataInNet comprises a data integration neural network and an auxiliary neural network. The data integration neural network is specifically designed to accept and integrate various types of data, including source terms, initial conditions, and boundary conditions. The auxiliary neural network is a fully connected network tasked with learning the residual physical information not captured by the data integration neural network. The forward propagation rule for data integration neural networks with $N^{*}$ modules are as follows:
@@ 4. Summary and discussion
Numerical experiments demonstrate the effectiveness of DataInNet. These experiments encompass the Helmholtz equation, the Poisson equation defined on an L-shaped domain, and PDEs with high frequency solutions. Future research will focus on devising more efficient data integration methods and assessing the applicability of this approach in more challenging scenarios.